### v3

# Jules Drean MIT CSAIL Miguel Gomez-Garcia MIT CSAIL Fisher Jepsen MIT CSAIL Thomas Bourgeat EPFL  and  Srinivas Devadas MIT CSAIL

Abstract.
Enclaves or Trusted Execution Environments are trusted-hardware primitives that make it possible to isolate and protect a sensitive program from an untrusted operating system. Unfortunately, almost all existing enclave platforms are vulnerable to microarchitectural side channels and transient execution attacks, and the one academic proposal that is not does not allow programs to interact with the outside world. We present Citadel, to our knowledge, the first enclave platform with microarchitectural isolation to run realistic secure programs on a speculative out-of-order multicore processor. We show how to leverage hardware/software co-design to enable shared memory between an enclave and an untrusted operating system while preventing speculative transmitters between the enclave and a potential adversary. We then evaluate our secure baseline and present further mechanisms to achieve reasonable performance for out-of-the-box programs. Our multicore processor runs on an FPGA and boots untrusted Linux from which users can securely launch and interact with enclaves. To demonstrate our platform capabilities, we run a private inference enclave that embed a small neural network trained on MNIST. A remote user can remotely attest the enclave integrity, perform key exchange and send encrypted input for secure evaluation. We open-source our end-to-end hardware and software infrastructure, hoping to spark more research and bridge the gap between conceptual proposals and FPGA prototypes.
1. Introduction
Secure Enclaves or Trusted Execution Environments (TEEs) (costan2016sanctum, ; aegis, ; ferraiuolo2017komodo, ; lee2020keystone, ; bahmani2021cure, ; feng2021scalable, ), are trusted-hardware primitives used to protect the execution of sensitive programs. They make it possible to run sensitive computations in isolation with strengthen privacy and integrity properties. Academic TEE proposals date back twenty years to XOM (xom, ) and AEGIS (aegis, ) and industry implementations include ARM TrustZone (trustzone04, ; pinto2019demystifying, ), Intel SGX (mckeen2013sgx, ; costan2016intel, ) and AMD SEV (sev2020strengthening, ).
Unfortunately, these kinds of platforms have been shown to be vulnerable to microarchitectural side channels and transient execution attacks (chen2019sgxpectre, ; koruyeh2018spectre, ; ryan2019hardware, ; van2020lvi, ). In particular, attacks such as those in (van2018foreshadow, ; van2020sgaxe, ) were able to not only extract sensitive data from Intel SGX enclaves, but also managed to break the security of the remote attestation mechanism.
One academic proposal (bourgeat2019mi6, ) has looked at securing enclaves in the context of microarchitectural side channels and transient execution attacks. They aim for a strong security guarantee where a potential attacker can only observe the enclave’s output. Unfortunately, this extreme isolation strongly restricts the class of programs that can be run, as no communication with the outside world is allowed except when entering or exiting the enclave.
Figure 1. Citadel’s Computing Stack
We are not the first to observe that restrictive enclave execution models limit the programmability and the performance of enclave programs(van2023cheri, ; yu2022elasticlave, ; li2021confidential, ; lee2022cerberus, ; park2020nested, ; steinegger2021servas, ) but those work ignore security challenges introduced by microarchitectural side-channels and transient execution attacks. When considering those challenges, even basic communication primitives like public shared memory with the operating system are not straightforward to implement efficiently and securely.
In this paper, we try to fill the gap between platform that have strong security but restrict enclave programs to the extreme and platforms that allow better enclave execution models but ignore side channels and transient execution attacks.
We first define our relaxed security goal: an attacker should only be able to observe the transcript of non-speculative accesses to a shared memory region. We then craft requirements for the hardware and the software so that, together, they can prevent secrets from leaking through speculative access to shared memory. In practice, the hardware will isolate private microarchitectural state and restrict the victim program speculation model to help it achieve the just described security guarantee. Intuitively, the more restrictive the speculation model, the lighter the requirements on the program, but the least performance we get.
In our case, we first consider a very conservative speculation model where no speculative shared accesses are allowed. This effectively prevents the existence of any speculative transmitters between the victim and the attacker. We describe how to enable shared memory securely for the specific open-source out-of-order design we used, while disabling speculative shared accesses. We also make sure to not enable any new speculative transmitters by bypassing the L1 D cache, and to not break address translation by tagging the TLB’s and the Address Translation Cache’s entries. Finally, we give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs on our FPGA prototype.
We use this secure baseline to evaluate the performance of the kind of programs that one could be interested in putting in an enclave. Out-of-the-box, we obtain an underwhelming $120.6\%$ performance overhead for a cryptographic library. We identify and analyze the performance bottlenecks to design a series of hardware-software mechanisms to improve both enclaves’ and unprotected processes’ performance. More precisely, we develop a reconfigurable Last Level Cache (LLC) partitioning scheme, a performance necessity for our small LLC. This reduces our overhead from $120.6\%$ to $19.4\%$. We also introduce software controls to enable a slightly less conservative enclave speculation model, where pipelined accesses to shared memory are allowed but predictors are hardwired to $pc+4$. The efficient way to use this mode is on code segments with burst accesses to shared memory. Under this less conservative speculation model, the programmer needs to make sure that straight-line speculation ($pc+4$) within these segments will not leak any secrets. This mechanism reduces overheads from $19.4\%$ to $2.7\%$ on our cryptographic benchmark.
Our final prototype is a modified out-of-order RISC-V 2-core processor synthesized to run on an AWS F1 FPGA. It fully implements all our security and dynamic resizing mechanisms, boots (untrusted) Linux and makes it possible for user-level applications to launch and safely communicate with secure enclaves executing real applications such as a lightweight Python runtime, a cryptographic library or a private inference application running a small neural network. We also implement an end-to-end attestation mechanism, making it possible for an enclave to identify itself independently of where the OS placed it in physical memory, enabling enclaves to act as a trusted third party. We implement and make all the pieces of software and modified hardware available, such as our secure bootloader, our security monitor and a Linux kernel module to interact with the security monitor. We summarize the different elements implemented for this paper in Figure 1. This makes our platform a unique environment for research in the domain of microarchitectural side channels, transient execution attacks and secure enclaves. We hope researchers will take ownership of the platform, attack it, modify it to prototype and test complex hardware-software mechanisms and research ideas.
To summarize, the main contributions of the paper are:
  • To our knowledge, we show the first end-to-end enclave platform running securely on an FPGA implementation of a speculative out-of-order processor.
  • We show how to build a shared memory mechanism while guaranteeing an attacker will only observe non-speculative accesses to shared memory.
  • Guided by early performance evaluation, we propose a series of hardware-software mechanisms to make the platform more configurable and improve both enclaves’ and unprotected processes’ performance.
We open-source our hardware, software and debugging infrastructure at https://github.com/mit-enclaves/citadel.
2. Motivation & Background
We want to : 1) run programs that can interact with the outside world. 2) formalize the guarantees that we expect the hardware to provide 3) have good performance.
2.1. Security Monitor
In many enclave platforms, including Citadel, the Security Monitor (SM) (costan2016sanctum, ) (lebedev2019sanctorum, ) is a small (9K LOC), trusted piece of software running at a higher privilege mode than the hypervisor or the OS. Its role is to link low-level invariants exposed by the hardware (e.g., if a specific memory access is authorized), invoking the right microarchitectural cleanup routine, and the high-level security policies defined by the platform (e.g., which enclave is currently executing on which core and which memory region does it own).
2.2. Microarchitectural Timing Side Channels
Microarchitectural side channels allow for information to be leaked from one security domain to another. These attacks exploit microarchitectural structures shared between different security domains, such as memory caches (gotzfried2017cache, ; brasser2017software, ), translation look-aside buffers (TLBs) (gras2018translation, ), branch predictors (evtyushkin2018branchscope, ; aciiccmez2006predicting, ), DRAM controllers (wang2014timing, ), etc., to transmit a signal. Because microarchitectural state cannot be directly observed through the Instruction Set Architecture (ISA), an attacker will typically rely on the timing of events to indirectly observe the microarchitectural state. For instance, if a piece of data has been loaded in the cache, accessing it will be faster than if it is only stored in main memory. Using various methods to prepare the shared microarchitectural state and observe the state changes due to the victim’s activity, the attacker can recover sensitive information about the victim’s execution, for example, which addresses have been accessed, which branches were taken in which directions, and/or information about victim program’s inputs. Cache side channels are the easiest channels to exploit, and so have been extensively studied, but any piece of shared microarchitectural state can theoretically be used as well.
To summarize the observations that could be made through side-channels, the community formalizes side-channels as a set of observations that an attacker could make about the microarchitectural state. For example, cache side-channels on unprotected machines typically lead to an observation model that includes the sequence of addresses of memory operations (sometimes with their timestamps) as part of the observation.
2.3. Transient Execution Attacks and Execution Models
Speculative or transient execution is one of the key performance features of modern processors. Unfortunately, it can make reasoning about security boundaries more difficult. Some microarchitectural data structures that orchestrate speculation (e.g., branch predictors) can be trained so the processor will execute code snippets speculatively, in ways that are incorrect or on data that should not be accessed during normal execution. Concretely, instructions that were not part of the program normal execution path can be executed out-of-thin-air, speculatively, by the processor. An example is the execution of code speculatively before the resolution of a preceding security check, such as a memory access bound. These erroneous executions happen speculatively, and these instructions will be squashed by the processor. Nonetheless, they will still change the microarchitectural state of the machine. For example, a speculative load to some data will not affect the values of the architectural registers but will load the data in the memory hierarchy and possibly change the state of caches. This means that a program that manages sensitive data can be coerced into leaking private data speculatively through a side channel. More importantly, once some specific code snippet, or gadget, has been identified, these attacks can be used to leak virtually any data located in the victim’s virtual address space. This type of attack, called Spectre, was first introduced in (Kocher2018spectre, ), and many variants have since been demonstrated.
3. Defining Security
To achieve security, we rely on both the hardware to enforce some isolation and on the fact the software verifies some requirements: we do not protect all-software out-of-the-box.
Intuitively, for a given security property, if the isolation enforced by the hardware is stronger, then the requirements on the software are weaker and we can enforce the security property for more programs.
Previous Work Previous work looking at enclaves in a context of transient execution attacks (bourgeat2019mi6, ) aims at limiting the attacker’s view to the enclave final output. Completion time was considered outside of the threat model and could be independently handled using traffic shaping techniques (padding) techniques. Unfortunately, this security property is too restrictive as it only considers a limited class of programs that do not communicate with the outside world.
Public Shared Memory We define a weaker security property in order to consider a more important class of programs to run securely. Our goal is for the system to guarantee that an attacker should only be able to observe the transcript of non-speculative accesses to a shared memory region. In particular, we don’t want an attacker to be able to observe any speculative accesses to shared memory from the victim nor any other microarchitectural state (accesses to private memory, branch predictors, TLBs etc…). However, similarly to previous work, we do not consider the timing of each access is observable by the attacker, i.e., there are no timestamps for the non-speculative shared memory accesses. We consider that leakage through the timing of shared memory accesses should be addressed using orthogonal traffic-shaping techniques such as (deutsch2022dagguise, ; wang2014timing, ; shafiee2015avoiding, ; zhou2017camouflage, ) and is left for future work.
Hardware & Software Requirements We now define what kind of requirements the hardware will needs to enforce and how the hardware restricts the victim program speculation model to help it achieve the just described security guarantee. 1) the hardware should properly isolate private microarchitectural state. 2) together, the hardware and the software need to prevent secrets from leaking through speculative accesses to shared memory. Intuitively, the more restrictive the speculation model of shared memory accesses enforced by the hardware, the lighter the conditions on the program to finally achieve security, but the least performance we get.
We first experiment with a restrictive speculative model where the victim program is not able to perform speculative accesses to shared memory. This effectively prevents the existence of any speculative transmitters between the victim and the attacker. The program only needs to make sure that the explicit shared memory accesses don’t leak information about the program secrets. Any constant-time cryptographic code satisfies these restrictions. Unfortunately this restrictive model gives us poor performance.
We experiment with a second speculative model called Burst mode where speculation on shared memory accesses is allowed, but all the predictors are deactivated and hard-coded to $pc+4$. The rationale is to allow pipelined accesses to shared memory to improve performance. Burst mode is intended to be only activated for code segments with straight-lined, burst accesses to shared memory. Under this less conservative speculation model, the software requirements are more important and the programs needs to verify that straight-line speculation ($pc+4$) within these segments will not leak any secrets. Outside of the said segments, we use the first speculation model.
3.1. Attacker Capabilities
We now complete our threat model by describing the attacker’s capabilities. The attacker is collocated on the same machine as the victim and can mount any software-based attack, including microarchitectural side channels and transient execution attacks. The attacker has compromised the majority of the software stack, including the operating system (OS) except for a limited trusted code base (TCB) including the secure bootloader, the SM (together 15K lines of code (LOC)) and the enclave. For the rest of this paper, we refer to any code outside of the TCB as the OS. We ignore attacks requiring physical access to the machine such as physical (e.g., noise, electromagnetic field, etc.) side channels. We also ignore side channels that can be mounted in software, but exploit the physical layer like power side channels. We do not consider Rowhammer attacks. As highlighted by our security definition, we are worried about enclave privacy and do not consider denial of service.
4. Citadel 1.0: A Secure Baseline
Even though defense mechanisms often propose to ”not speculate” when executing some instructions that can dirty the microarchitectural state, implementing such policies is not completely straightforward. In this section, we describe how to securely enable shared memory for our specific out-of-order open-source design while disabling speculative shared accesses. We also make sure to not enable any new speculative transmitters by bypassing the L1 D cache, and to not break address translation by tagging the TLB’s and the Address Translation Cache’s entries. Finally, we give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs.
4.1. Hardware Isolation
Memory Isolation Sanctum (costan2016sanctum, ) already presents hardware mechanisms for memory isolation in the context of an untrusted OS. Like RISC-V Physical Memory Protection (PMP) (waterman2015risc, ), Sanctum introduces a mechanism to partition physical memory and prevent a malicious OS from accessing another security domain’s private memory regions, or different enclaves from accessing each other’s private memory. At its core, the mechanism uses two 64-bit registers per core (one for trusted memory and one for untrusted/shared memory) as memory bitmaps. Each bit represents a fixed-size physical memory region (typically 32MB) and functions as an access-rights flag for the currently-running program. Citadel borrows this mechanism for its simplicity and enforces these checks on every memory access, including speculative accesses and for accesses performed during page walks.
Microarchitectural Isolation Following MI6 (bourgeat2019mi6, ) methodology we partition all spatially shared microarchitectural state (LLC, MSHRs…) and flush temporally shared state on context switch (all cores’ pipeline state, L1 caches, TLBs…). We are not vulnerable the DRAM controller side channels as we inherit MI6’s constant time DRAM padding mechanism. Finally, our processor does not support hyper-threading and we are not vulnerable to related side channels such as port contention.
4.2. Enabling Shared Memory
Figure 2. The Dual Page Table Mechanism
Sanctum also uses a dual-page-table mechanism (see Figure 2) to differentiate between trusted and untrusted memory based on virtual addresses, which we also borrow in Citadel. Two new registers (equivalent to base and bounds) are used to define the trusted virtual memory range (i.e. the enclave’s virtual memory). The SM enforces the invariant that the physical memory mapped as trusted is owned by the corresponding enclave. When a memory access to a virtual address is performed by a program, one of the two page tables needs to be picked to translate the virtual address. If the virtual address falls within the trusted range, the translation is performed using the enclave page table. Another important invariant enforced by the SM is that the enclave page table is always stored within enclave private memory. If the virtual address falls outside the trusted range, it corresponds to an untrusted or shared memory access and the translation is performed using the OS page table. The OS page table is untrusted and stored in OS memory. Note that for every physical address used during address translation, additional bound checks are performed to ensure they belong to the same security domain (enclave or OS memory) as the translated address. This mechanism is especially handy to eliminate speculative transmitters, as there is no need to perform any speculative address translation (which could leak information) to know whether a memory access is trusted or not.
4.3. Restricting Speculative Access to Shared Memory
In the baseline out-of-order processor we are hardening, memory instructions are decoded, renamed and then simultaneously queued in the reorder buffer (ROB) and the memory reservation station (MemRS). When an instruction’s operands are ready, it is dispatched to the memory execution pipeline (MemExePipeline). Our approach identifies instructions that access untrusted memory early in the MemExePipeline and delay the translation and processing of the memory accesses until the instruction is no longer speculative, i.e., when it reaches the head of the ROB. These changes require the modification of MemRS and MemExePipeline.
Figure 3. Changes to Memory Execution Pipeline. New elements are shown in blue.
Modifying the Memory Execution Pipeline Upon entry into the MemExePipeline, checks are performed on memory instructions to decide if the memory operation is safe to translate and then issue, or not. It is important to perform these checks very early as the next steps in the MemExePipeline touch shared microarchitecture (through address translation and cache hierarchy for instance) potentially transmitting information to an attacker. A memory access is deemed safe if it accesses trusted memory OR if it is non-speculative. Because the enclave trusted memory range is defined in virtual memory, this check can be performed as soon as operands (i.e. virtual address) are available. If an instruction tries to access untrusted or shared memory, we check that it is non-speculative by conservatively checking if it has reached the head of the ROB. If the instruction is safe, it is dispatched to the next stage of the MemExePipeline. If it is not safe, the instruction is squashed and a signal is sent to the MemRS indicating that it should re-dispatch it once it reaches the head of the ROB.
Modifying the Memory Reservation Station Our mechanism requires MemRS being able to re-dispatch some instructions once they become non-speculative: untrusted memory loads will indeed typically need to get dispatched twice.
Disabling the Mechanism for the OS Blocking speculation on memory access can have a significant performance impact. The SM makes sure that nonenclave programs (typically the OS and standard usermode programs) are still allowed to speculate for their memory accesses by disabling the trusted memory mechanism.
4.4. Speculative Transmitters in the L1 D Cache
The core-private part of the memory system, including the L1 D cache, is now shared between trusted memory accesses and untrusted or shared memory accesses. While shared memory cannot be accessed speculatively, trusted memory can. It means that shared memory can be speculatively evicted by a speculative trusted memory access. Furthermore, this eviction will be observable to an attacker. The cache coherency protocol is such that writing to a memory location that is cached by another core will take more time than writing to a memory location that is not cached by another core. As a result, a shared L1 D can be leveraged by an attacker to create a speculative transmitter and mount a spectre-style attack. To our knowledge, we are the first to identify this subtle attack vector. Note that this doesn’t only apply to shared cache sets, but it also applies to contention over other shared resources in the L1 D cache (including the pipeline to access the SRAM). Partitioning the L1 D cache would be too costly, instead we bypass the L1 D cache for untrusted or shared accesses and directly send the requests to the LLC.
4.5. Tagged TLB for Correct Address Translations
When using the dual page table mechanism, the core-private part of the memory system is shared between trusted memory accesses and untrusted memory accesses. For functional correctness, we need to make sure that addresses are translated correctly for each domain. This small microarchitectural structure accelerates address translation by caching intermediate results used by the page walker. Unfortunately, only a few upper bits of the virtual address are used at the beginning of a page walk, easily resulting in aliasing across the two domains. We tag the Address Translation Cache by adding a bit to each entry indicating if the cached information applies to the trusted or untrusted memory address translations. Aliasing can also occur in the TLB if the trusted virtual memory range is smaller than the granularity of a virtual memory page (which can be if we use Mega or Giga Pages). We also tag the TLB entries to avoid address mistranslations.
4.6. Miscellaneous
We statically partition MSHRs between two cores and place a fair round-robin arbiter at the entry of the LLC. Because these two mechanisms only marginally affect the latency of the LLC for a 2-core design, we observe negligible ($\approx$ 0.1%) performance impact.
We also expose controls for speculation and predictors to the software through Control Status Registers (CSRs). Specifically, a program is able to deactivate speculation all together (every instruction waits to reach the head of the ROB), delay the issuing of shared memory accesses to the head of the ROB, deactivate the training of predictors (branch predictors and return address stack) and/or their use (and then it falls back to the predictor $pc+4$).
4.7. Building the Software Infrastructure
We are interested in evaluating our platform on a Linux-ready system and for programs that one would be interested in putting in an enclave (e.g. cryptographic libraries). This means a lot of software needs to be developed to be able to launch enclaves from a Linux user mode process.
First we implement the SM at the most privilege level of RISC-V using a mix of assembler and C and add the necessary pieces of a micro-kernel required to boot Linux on top of it. We also implement what we call a Mini-SM which is a downsized copy of the SM private to each enclave. This avoids contention over most SM resources between enclaves and untrusted programs. To protect the SM against transient execution attacks, we disable speculation during SM execution. Second, we develop a Linux Kernel module that allows user programs to request SM calls and set up enclaves, keeping Linux responsible for the resource management. Finally, we implement a Secure Bootloader and an end-to-end attestation mechanism that makes it possible to uniquely identify an enclave, establish a shared secret with it and exchange encrypted private data. Our attestation mechanism makes it possible for an enclave to be placed anywhere in physical memory and still have consistent identity through the measurement of the direct mapping between virtual addresses and memory content without measuring physical addresses. The unique mapping is attested and enforced by the SM during enclave setup. More implementation details can be found in the evaluation section (see section 6).
5. Citadel 2.0: Improving Performance
Our first baseline successfully achieves security. In this section, we show several mechanisms to improve its performance. First. fixed LLC partitioning á la MI6 destroys performance for our 2-core machine. We show a reconfigurable LLC partitioning scheme and a fine grain flushing mechanism. Furthermore, disabling speculative shared memory accesses makes software requirements especially light, but provides limited performances. We show a relaxed speculative model where we enable pipelined accesses to shared memory but disable predictors and hard-code them to $pc+4$. When this mode is enabled, the program needs to verify more requirements, i.e., that no secret is leaked when speculating $pc+4$. We show that this new mode improves performance for code segments with burst shared memory accesses, but it hampers performances in the general case. We observe that this is due to the TLB being trashed by misspeculations across functions context. This is an effect that was studied (in a nonsecurity context) in the early 2000s(onur1, ; onur2, ; onur3, ): misspeculations can induce large overhead not only due to the delay in executing the right path, large delays are induced due to the trashing of the TLBs. Finally, we show how insights from building the SM make it possible to securely optimize context switch and enable frequent interrupts.
5.1. Evaluation-Guided Optimisations
Our secure baseline runs on FPGA with a Linux environment (see section 6 for our full configuration) which makes it possible to analyses performance bottlenecks on real programs. For this section, we use an elliptic curves (Ed25519) cryptographic library as a running example (see section 6 for more details). At a high level, our enclave interactively receives messages to sign from a benchtest program using shared memory. Out-of-the box, we see an underwhelming overhead of 120% overhead compared to directly (and insecurely) linking the cryptographic library to our benchtest program.
5.2. Reconfigurable LLC Partitioning
An obvious bottleneck in our design is cache partitioning. Sanctum and MI6’s mechanism (i.e. dividing the LLC in 64 equal size slices to match the 64 DRAM regions) is a terrible idea for our system. Our 2-core prototype has a 1MB LLC (see 5), which divided by 64 gives us 16KB LLC slices. This is smaller than the 64KB L1 D+I caches. Because our caches are inclusive, this means an enclave or a program that resides in one memory region cannot even fully use the L1 caches. MI6 circumvented that issue by assuming a 16-core design with a large 16MB LLC.
We build a mechanism for the OS to reconfigure the LLC partitions to improve enclave and other program’s performance. This costly operation is only required when changing the LLC layout, which only happens before launching an application or an enclave with an important memory footprint. The OS requests a reconfiguration through an SM call while no enclaves are running. The SM will first check that the new partitioning is valid, doesn’t affect regions assigned to the SM. It then waits for other cores to flush their L1 caches and to enter a blocked state in the SM, making sure no requests will be sent to the modified regions during reconfiguration (which could put the cache in an incoherent state). The SM will then interact with the hardware to flush the LLC slices that are modified and set the LLC in its new configuration.
At the hardware level, we now design a practical reconfigurable LLC partitioning scheme. We cannot use way partitioning (kiriansky2018dawg, ) as the LLC has only 16 ways, and we need up to 64 slices, and Ironhide’s (ironhide, ) strategy assumes limited reconfiguration. We propose a new, simple, configurable set partitioning scheme that only requires minimal changes in the hardware.
The configuration for each LLC slice (base and bound of the set-index range) is stored in a big register (1216 bits) called the set-index range table. The SM can modify the entries of the set-index range table by sending MMIO requests to the cache, which is seen as a new device. When the LLC receives a request, it will compute the new set index for the corresponding address. We represent the transformation used in Figure 4. First, we extract a region ID from the address upper bits and obtain the corresponding index range base and size by looking up the set-index range table 1. We then extract the address original set index, located in the middle of the address, and tweak it. We compute the original set-index value modulo the size of the index range 2 and add the base index 3. This effectively restricts the new set index to the corresponding index range, partitioning the LLC 4. Note that we must extend the length of the tags used by the LLC to include the entire original index in case an index range is of size 1 and all addresses from the same region map to the same set.
We observe on our cryptographic library that we reduce overheads from 120% to 19.4% by using a non-uniform partition for the LLC. More details can be found in Section 6.5.
Figure 4. LLC set indexing scheme. The function takes an address as an input and output a set index in the LLC slice allocated to its memory region.
LLC fine-grain flush To go hand-in-hand with our reconfigurable LLC partitioning, we also want to be able to selectively flush parts of the LLC. We create a Zero Device in the upper physical address space. It is implemented as a simple piece of logic connected to the cache hierarchy at the same level as the DRAM, but that always sends back zeros. The trick is that the address space covered by the zero device has the same size as the DRAM and is located at addresses that will collide with DRAM addresses in the cache. Once the SM has placed the machine in a correct waiting state (where all the cores have flushed their L1 and are waiting in the SM), it only needs to access an eviction set in the zero-device address space to flush the cache slice. This means that flushing a small LLC slice is much faster than flushing a big one. We look at the performance of our flushing mechanism in the evaluation section (see 6.5).
5.3. Pipelined Shared Accesses
At this step, we observe that our test benchmark still present a 19.4% overhead compared to our insecure benchmark. We decide to measure the performance of different sections of the code to find and analyze the bottleneck. We identify the function creating the slowdown to be located inside the SHA256 library. This function performs burst accesses to shared memory to load the message to be signed. This behavior is exacerbated by the compiler, trying to optimize the code by placing all the load instructions to shared memory one after the other. Indeed, with our current mechanism, memory accesses to shared regions, first cannot use the L1 and have a 10 cycle latency, second cannot be pipelined as each instruction needs to wait to reach the head of the ROB to be sent. This is especially costly during burst accesses. The LLC supporting 16 outstanding requests and having a 10 cycles latency, 16 memory accesses should be pipelined to take a total of $10+16=26$ cycles. Without pipelining, they will now take $10*16=160$ cycles.
As a result, we designed a mode where we temporarily change the speculation model to allow shared memory burst accesses (as pipelining is a form of speculation). Nevertheless, we don’t want to do so while enabling any type of speculation. Indeed, we don’t want our program to speculatively jump to a spectre-like gadget and transmit secrets to our attacker. To avoid this, we tweak the speculation model to allow pipelined accesses to shared memory but disable the use and training of any predictors that could result in the speculative execution of rogue code. Specifically, we disable the Branch Target Buffer (BTB), the Branch History Table (BHT) and the Return Address Stack (RAS). Effectively, that guarantees that the hardware will always predict that the next program counter to be “$pc+4$” and avoid speculative execution of arbitrary code. We call this straight-line speculation.
Modifying the speculative model also modifies the hardware software contract. The programmer now needs to ensure that when this mode is enabled, no speculative transmitter can be reached by the program when doing straight-line speculation. Static analysis can be done manually or by reusing many existing tools: e.g. oo7(wang2019oo7, ), Kasper(johannesmeyer2022kasper, ), Kleespectre(wang2020kleespectre, ), Spectector(guarnieri2020spectector, ). Note that the usual critique and limitation of these tools are that 1) the code bases are too big to analyze without producing a lot of false positives 2) they make simplifying assumptions on the speculation model. Interestingly, both of those limitations are not concerns in the context of Citadel’s burst mode because the code snippets we analyze are small (e.g. boiling down to a memcpy function for instance) and the speculation model is well-defined, simple and enforced (only speculates PC+4). Hence, these works are well-suited to our needs. We consider adapting these tools to our more simple setup as future work. We add two new configurations in our software-exposed configuration registers to enable and disable this “burst” mode. The right way to use this mode is, like in our case, during burst accesses to shared memory (e.g. a memcpy function moving data from public memory to private enclave memory). The programmer only needs to identify these series of burst access and guard them with instructions activating and deactivating burst mode. When doing so on our cryptographic benchmark, we observe that overhead falls from $19.4\%$ to $2.7\%$. More results are presented in the evaluation section.
5.4. Relaxed Flushing Policies
Because we build the entire software infrastructure required to run enclaves, we were able to observe that some security domains, especially the enclave and some parts of the mini-SM, do not have secrets to keep from each other. Specifically, the mini-SM never manipulates sensitive data. As a result, the mini-SM does not need to flush microarchitectural state when context switching to enclaves. This reduces the cost of mini-SM call and opens the door to the use of frequent software or timer interrupts. To disable flushing selectively, we modify the hardware to deactivate systematic flushing and exposed to the SM a register to trigger the flush. We write the SM to trigger the flushing mechanism on required context switches, but not on most traps and SM-calls from the enclave. We evaluate the performance implication of our mechanism in Section 6.6.
6. Evaluation
6.1. Experimental Setup
Frontend
2-way superscalar, 256 entries BTB, 8 entries RAS, tournament predictor
Reorder Buffer
64 entries, 2-way insert/commit
Reservation Stations
2 stations for ALU (16 entries), 1 for FPU (16 entry), 1 for memory (16 entries)
L1 (I/D)
32KB, 8 ways, 8 outstanding requests
Load-Store Unit
24 LdQ entries, 14 StQ entries, 4 SB entries
LLC (L2)
1MB, 10 cycles latency, 16 ways, max 16 outstanding requests
Memory
120 cycles latency (24 outstanding requests)
Figure 5. Citadel’s microarchitectural configuration.
Citadel’s processor is a modified out-of-order 2-core based on Riscy-OOO (zhang2018composable, ) and MI6 (bourgeat2019mi6, ). The parameters used for the cores and the memory subsystem are given in Figure 5. Citadel runs on an AWS EC2 F1 instance, equipped with a Virtex UltraScale+ FPGA. On this FPGA board, Citadel can be clocked at $30$ MHz, which is similar to the baseline design. We use the RISC-V tool chain with GCC 12.2.0. We compile all software with the -O3 flag. We run Linux Kernel 6.2.0. Unless mentioned differently, all performance numbers are collected by running the benchmark on the FPGA using performance counters and an average of ten runs.
6.2. Implementation Data and TCB Breakdown
Component
LOC Size
Processor (bourgeat2019mi6, ) (+ modification)
$\sim 60$ KLOC (+ 250 ) N/A
Bootloader
5236 55KB
Security Monitor
9287 97KB
Mini-SM
2301 5.6KB
Linux
$\sim 20$ MLOC 157MB
SM Kernel Mod.
507 505KB
TCB Software
14523 152KB
Table 1. Software Component Size and TCB Breakdown. Mini-SM is a subset of the SM code. Lines in red correspond to elements excluded from the TCB.
Number of lines of code (LOC) and binary size of Citadel’s components are presented in Table 1. Due to their limited functions, the SM and the Bootloader can be kept small and the total software Trusted Code Base (TCB) is less than 15K LOC. This helps to minimize potential bugs and reduce the software attack surface. It can also help to make future formal verification efforts tractable. Final TCB size needs to include the enclave’s binary, which depends on the application.
6.3. Evaluating the End-to-End Platform
MicroPython Runtime We do a minimal port of MicroPython (microphython, ) inside one of our enclaves. We start with the bare-metal ARM implementation. We modify the configuration files and the linker script so our binary can run in the virtual memory layout provided by our enclave. We only need to implement a few functions by hand. For instance, we redirect the console to print to shared memory and need to write a bit of assembly for the setjmp and longjmp functions used for exception handling. We also add a cryptographic library, so the enclave can sign the console output using its own private keys recovered through the remote attestation mechanism. A hash of the standard output is extended on each character print and signed on the enclave exit. We cross-compile MicroPython using the RISC-V tool-chain, build without standard library, and generate position independent code. In total, we need to manually modify 128 LOC for the port and add the 5K LOC cryptographic library. The final compiled binary measures 180KB. With the enclave’s code added, the TCB is now 65,174 LOC, most of it (70%) being MicroPython.
Figure 6. MicroPython Enclave Execution Breakdown.
We use a simple primality test as a benchmark where a number is found to be prime if none of the numbers smaller than it can divide it. We test the primality of an arbitrary small prime number: $19,531$. The tests take an average of $4.1\times 10^{7}$ cycles to execute, with 51% of the time spent within the Python program. This means the enclave and attestation mechanism add 94% overhead for this small computation. We show a visual breakdown of the running time in Figure 6. We note that most of this overhead comes from hashing enclave code (25% of the total running time) but also generating the enclave keys and signing the certificate (10% of the running time). These two numbers are especially high since this code is run non-speculatively inside the SM. On another hand, hashing and signing the output inside of enclave takes a negligible amount of time (less than 1% of the total running time). Because this cost is fixed, it could definitely be amortized for bigger applications.
Securing a Cryptographic Library We write a wrapper around an ED25519 RISC-V library (ed25519github, ). The wrapper makes it possible to keep the keys used by the library inside of the enclave and for the library functions to be accessed in a Remote Procedure Call style (RPC). We add a queue to shared memory, so untrusted applications can send requests to the library to generate keys (that will stay in enclave memory), or to sign messages using previously generated keys. In total, we need to implement 334 LOC of C, which is to be compared to the 4,514 LOC of the crypto library (7% overhead). With the enclave’s code added, the total TCB is now 19,231 LOC. The final crypto-enclave binary is 71KB.
We compare performance between three scenarios. In each of these scenarios, we measure the number of cycles required to sign randomly generated messages. We split 1MB of randomly generated data into $\approx 3000$ packets of various lengths following the arbitrary Internet Mix distribution (imix, ). In the first setup, which corresponds to our insecure baseline, we do not use any enclave as we directly link the cryptographic library to our user-level application. In the second scenario, called “multi-core”, the enclave constantly runs on a dedicated core and programs running on other cores can forward requests through the shared memory mechanism. In the third scenario, termed “re-entry”, the enclave is re-entered every single time an application calls the cryptographic library on the same core as that of the application. We see very little difference (2.9%) between the first two setups, and a more important 27.6% slowdown between the first and third scenario. Indeed, our cryptographic algorithm is compute-heavy, making the communication overhead negligible in the second scenario. Nevertheless, context-switching in and out of an enclave is expensive as it requires flushing the microarchitectural state, which makes the slowdown more important in the re-entry scenario. These two secure solutions highlight an interesting tradeoff between resource occupation (one core and a DRAM region vs. only a DRAM region) and performance.
Private Inference We evaluate the following scenario where a remote user wants to do perform some private inference task on a neural network evaluated on an untrusted server. We pace a simple neural network model inside of one of our enclaves. We use onnx2c (onnx2cgithub, ) to convert a ONNX file describing a handwritten digit recognition model (trained on MNIST) to standard C code and run private inference inside of our enclave. The user uses our remote attestation mechanism to verify the enclave identity and establish a shared secret with it. The user is then able to send 1024 encrypted requests to the enclave that are passed along to the enclave by the untrusted OS. Upon reception, the enclave decrypts the ciphertext and performs the inference task on the private input. It then sends back an encrypted output to the remote user. In total, we need to implement 227 lines of straightforward C, which is to be compared to the 1536 LOC of the neural network and the 410 LOC of the AES library and the 4386 LOC of ED25519 we now need to include inside of the enclave for encryption and key exchange. With the enclave code added, the TCB is now 21K LOC and the final enclave binary measures 54KB.
We now compare the performance between our setup and an insecure baseline where the model is directly evaluated outside of an enclave and where the LLC is not partitioned. We do consider that incoming traffic from the remote user is encrypted and needs to be decrypted first. In both setups, runtime is split 55% and 45% between AES operations and inference. We see negligible (0.76%) overhead for running a model inside of the enclave, with all of it coming from the necessity to copy the ciphertext in enclave private memory before performing decryption. This is not surprising as the rest of the computation happens in enclave private memory and should not suffer from any slowdown. This doesn’t account for the time required to set up the enclave, which is a one-time cost but represents 23.3% of the cost of performing 1024 inference.
6.4. Evaluating Secure Shared Memory
We first demonstrate that naive public memory sharing is vulnerable to transient execution attacks by mounting an attack on an insecure version of Citadel where speculation on shared memory is allowed. The attack uses the LLC as a side channel and use a specter gadget located in enclave code to leak arbitrary enclave memory to untrusted code ruining on another core. Anecdotally, we were not able to successfully mount the attack on our secure prototype.
Table 2. Normalized runtime for signing and hashing three thousand random messages sent by the OS to an enclave. Runtime is normalized to the insecure baseline where the library is directly linked to the testbench program. We compare Citadel with two other configurations where the library code is not annotated for burst accesses to shared memory (- Burst) or all annotated (++ Burst).
[…]
We then evaluate the perfromance overhead of our comunication mechanism. We use a similar setup as Section 6 but only measure time spent computing the cryptographic function. The enclave performs $3000$ cryptographic operations (signature or cryptographic hash) over random inputs provided by the OS. We compare three setups. First, our final design with burst mode enabled around shared memory burst sequences (Citadel). Then Citadel without annotating code for shared memory burst sequences (- Burst). Finally, another (insecure) configuration where we annotate all the code to be run in burst mode (++ Burst). Our results are presented in Table 6.4. We report the runtime normalized to the insecure baseline (directly linked library and no LLC partitioning). We observe an overhead of 2.7% on the signature benchmark and 46.6% for the hashing benchmark. This is because the hashing application accesses shared memory much more frequently compared to the signing application that is more computationally bounded. The hashing function is more affected by the loss of the L1 cache. This difference between the benchmarks is also highlighted by our other results, where not annotating the code for burst shared accesses has an $397.3\%$ overhead for the hash application and a $19.4\%$ for the signing applicaiton. These results also demonstrate the efficiency of our burst mechanism when used properly.
We also investigate the performance of using this mode on the whole enclave code instead of the limited code segment when the enclave communicates. On our benchmarks, we observe that the performance overhead goes up to $65.4\%$ for the signing benchmark and $395.9\%$ for the hashing benchmark. We investigate the origin of this overhead and realize that miss speculation due to the deactivation of the RAS generates a lot of random load that pollutes the TLB, triggers many page walks and terribly slows down the virtual memory subsystem. This is due to many small functions being linked one after the other in the binary. Some of these functions start with loads using generic registers as addresses (for instance when accessing a pointer passed as an argument to the function). When encountering a ret instruction, the processor will first speculate that the following instructions should be executed. In this speculative window, it will start running the instructions of the function linked right after and perform loads or stores to random addresses from registers containing garbage. We confirm this hypothesis by instrumenting the binary and observing that when the RAS is reactivated (but not other predictors), the number of TLB misses, page walks and the overall execution time go down significantly. As a result, we recommend a hybrid approach where the burst mode is only used when the enclave wants to communicate.
Table 2. Normalized runtime for signing and hashing three thousand random messages sent by the OS to an enclave. Runtime is normalized to the insecure baseline where the library is directly linked to the testbench program. We compare Citadel with two other configurations where the library code is not annotated for burst accesses to shared memory (- Burst) or all annotated (++ Burst).